# New Advances in Phonons: From Band Topology to Quasiparticle Chirality

### v1

@@ IV.1.3 Body-centered Silicon: $\mathbb{Z}_{2}$ Dirac Phonon
Figure 27: $\mathbb{Z}_{2}$ Dirac phonon in body-centered silicon with space group $Ia\bar{3}$. (a) Crystal structure of silicon. (b) Bulk Brillouin zone (BZ) and (001) surface BZ. (c) Phonon band structure, where the colored dots denote two distinct $\mathbb{Z}_{2}$ Dirac phonons, $P_{1}P_{3}$ and $P_{2}P_{2}$, characterized by different irreducible representations (irreps). (d)-(e) Surface states and Fermi arcs associated with the $\mathbb{Z}_{2}$ Dirac phonons of irreps $P_{1}P_{3}$. (f)-(g) Surface states and Fermi arcs corresponding to the $\mathbb{Z}_{2}$ Dirac phonons of irreps $P_{2}P_{2}$. Adapted from Ref. Zhang and Murakami (2023).
As discussed in Sec. II.2, conventional Dirac phonons are topologically trivial because they consist of two Weyl phonons with opposite Chern numbers, which do not guarantee gapless helicoid surface states. To provide an intuitive understanding of $\mathbb{Z}_{2}$ Dirac phonons and their characteristic anti-parallel helical surface states, we will take an experimentally synthesized material candidate as an example in this subsection, showing their topological properties Zhang and Murakami (2023).
@@
[…]
Similarly, Figs. 27 (f) and (g) show the surface states and arcs for the $P_{2}P_{2}$ Dirac phonon, which also exhibit two pairs of anti-parallel helical surface states around $\bar{M}$. Thus, both the $P_{2}P_{2}$ and $P_{1}P_{3}$ Dirac phonons in silicon’s phonon spectrum are $\mathbb{Z}_{2}$ Dirac phonons, each hosting two pairs of anti-parallel helical surface states (termed anti-parallel quad-helical surface states in Zhang and Murakami (2023)). The (anti-)parallel quad-helical surface states represent the maximal number of surface states achievable for (Dirac) Weyl phonons in solids, as constrained by onsite and crystalline symmetries.
Figure 28: (a) Conventional and primitive unit cells of MoB2. (b) Phonon spectra of MoB2. (c) Topological nodal line identification in MoB2, determined through compatibility relations and symmetry-based indicator analysis. (d)-(e) Side view and [111]-oriented top view of the nodal lines near 30 meV in MoB2, exhibiting a quantized $\pi$ Berry phase. (f) Bulk Brillouin zone and [010]-projected surface BZ of MoB2. (010) surface local density of states (LDOS), showing both bulk-derived modes and surface-unique modes (left panel: integrated intensity). Drumhead surface phonon states (within white dashed rectangles) connecting nodal line crossings display maximal LDOS, as evidenced by sharp intensity peaks at nodal-line energies (shaded regions). Adapted from Ref. Zhang et al. (2019b).
@@ IV.2 Topological Phonons with 1D Degeneracies
For MoB2, our symmetry analysis yields the indicator $z_{2,1}z_{2,2}z_{2,3}z_{2,4}z_{4}=(1110)$. Following the theory outlined in Sec. II.3 (Eqs. 72 and 73Song et al. (2018); Zhang et al. (2020a), this indicator predicts:
@@
To demonstrate the topological surface states associated with these $\mathcal{PT}$-protected nodal lines, we present the local density of states (LDOS) along the (010) direction in Fig. 28 (f), corresponding to the surface Brillouin zone shown in Fig. 28 (e). Unlike Weyl points that exhibit helicoid surface modes Fang et al. (2016); Zhang et al. (2018a); Miao et al. (2018); Zhang et al. (2020b, 2022b); Zhang and Murakami (2023), the nodal lines generate characteristic drumhead-shaped surface states, as indicated by the white dashed box. Notably, while the bulk phonon density of states shows minima near $\sim$7.25 THz and $\sim$23 THz, the surface modes display a pronounced peak in LDOS at these energies (Fig. 28 b). The flat surface states could induce electronic anomalies through electron-phonon coupling, potentially leading to unique surface phenomena.
@@
In 2D systems, Eq. (75) typically defines isolated $\bm{k}$-points, corresponding to 2D Dirac semimetals Li et al. (2020a); Jin et al. (2018b). In 3D systems, however, the same condition can be satisfied along closed $\bm{k}$-space contours, giving rise to topological phononic nodal lines. Numerous material candidates hosting such nodal lines have been theoretically predicted and experimentally observed. Based on their geometric configurations, these nodal lines can be categorized into several distinct types: helical nodal lines Zhang et al. (2019b), straight nodal lines Li et al. (2020b); Liu et al. (2021a), nodal rings Jin et al. (2018a), nodal chains Chen et al. (2021b), and nodal net Liu et al. (2021c). Phononic topological nodal lines exist ubiquitously in centrosymmetric solids Liu et al. (2022).
Straight nodal lines along high-symmetry lines (HSLs) with $C_{n}$ rotational symmetry can be categorized into two fundamental types based on their transverse dispersion: linear and quadratic types Liu et al. (2021c, 2022); Qin et al. (2024). This classification applies particularly to systems with threefold rotational symmetry ($n=3$), where the low-energy effective Hamiltonian takes the general form:
@@
where $k_{\pm}=k_{1}\pm ik_{2}$ with $k_{1}$ and $k_{2}$ being the wave vector perpendicular to the rotational axis. $A$ and $B$ are complex constants that do not depend on $k_{1}$ and $k_{2}$. The location of nodal line is determined by $Ak_{-}+Bk^{2}_{+}=0$ which give rise to 4 nodal lines: One located at the rotation axis along $k_{3}$ with Berry phase $+\pi$, while the other three located at generic positions which are related by $C_{3}$ rotation and have Berry phase $-\pi$, i.e.,
$\begin{split}(0,0,k_{3}):&\textrm{Berry phase}\leavevmode\nobreak\ +\pi\\ \left(k_{0}\cos\frac{\theta_{0}}{3},k_{0}\sin\frac{\theta_{0}}{3},k_{3}\right):&\textrm{Berry phase}\leavevmode\nobreak\ -\pi\\ \left(k_{0}\cos\frac{\theta_{0}+2\pi}{3},k_{0}\sin\frac{\theta_{0}+2\pi}{3},k_{3}\right):&\textrm{Berry phase}\leavevmode\nobreak\ -\pi\\ \left(k_{0}\cos\frac{\theta_{0}-2\pi}{3},k_{0}\sin\frac{\theta_{0}-2\pi}{3},k_{3}\right):&\textrm{Berry phase}\leavevmode\nobreak\ -\pi\end{split}$ (77)
where $k_{0}=|-\frac{A}{B}|$ and $\theta_{0}=\mathrm{arg}\left(-\frac{A}{B}\right)$. Nodal lines can be classified into two fundamental types based on their transverse dispersion: (i) linear nodal lines with $\pi$ Berry phase (topologically nontrivial) and (ii) quadratic nodal lines with $2\pi$ Berry phase Liu et al. (2022); Qin et al. (2024). The topological character is determined by the dominant term in the effective Hamiltonian: when the linear term dominates ($|A/B|\gg 1$), the system exhibits a $\pi$ Berry phase, whereas dominance of the quadratic term ($|A/B|\ll 1$) yields a $-2\pi$ Berry phase. For systems with $C_{4}$ or $C_{6}$ symmetry, the linear term in Eq. (76) vanishes, allowing only a single straight nodal line along the rotational axis. In these cases, quadratic nodal lines emerge with Berry phase ±$\pm 2\pi$, corresponding to a winding number $n_{W}=\pm 2$ Liu et al. (2022). This distinct topological signature arises from the doubled winding of the pseudospin texture around the nodal line.
Figure 29: Two types of unconventionality. Atomic vibrations act as $p$ orbitals. (a) Phonon modes contributed by the vibration center at an empty site. (b) Phonon modes contributed by the vibration center at an atom site, yet show a non-atomic vibration orbitals (non-$p$ orbital symmetry). (c) Phonon spectra of the monolayer black phosphorous. (d) Phonon spectra for the monolayer black phosphorous with an open boundary condition, where there are floating obstructed surface phonon modes $\sim$ 9 THz. The real space vibration for the surface phonon mode marked by the green hexagon is shown in (e). Adapted from Ref. Zhang et al. (2023a).
@@ IV.3 Obstructed Atomic Insulator-like Topological Phonons
Phonons represent quantized collective vibrations of atomic lattices, where each mode corresponds to a distinct pattern of atomic displacements. However, certain phonon modes exhibit unconventional origins: they may arise from vibrations localized at empty atomic sites or involve non-atomic orbital symmetries (e.g., non-$p$-orbital-like vibrations), as illustrated in Figs. 29 (a)–(b). These anomalous modes, termed obstructed phonon modes Zhang et al. (2023a, b); Ma et al. (2023); Xu et al. (2024b), defy conventional descriptions of lattice dynamics. The unconventional nature of these modes can be identified through either real-space invariants or open-boundary calculations Zhang et al. (2023a, b). The former one quantifies the mismatch between atomic positions and vibrational centers, while the latter one reveals localized surface modes (“floating states”) detached from bulk phonon spectra. Such modes typically emerge from dangling-bond vibrations on crystal surfaces and are intimately linked to the material’s symmetry and electronic structure. Their spectral isolation and spatial localization make them promising for controlling phonon-mediated phenomena at interfaces or defects.
As a concrete example, Figs. 29 (c) and (d) display the phonon spectra of monolayer black phosphorus under periodic and open boundary conditions, respectively Zhang et al. (2023a). Notably, the open-boundary spectrum reveals floating obstructed surface phonon modes near 9 THz, which are absent in the periodic calculation. The real-space vibrational pattern of one such surface mode (indicated by the green hexagon in Fig. 29 (d) is shown in Fig. 29 (e), clearly demonstrating its localized character at the edge.
Figure 30: (a) First-principles calculations of right-handed tellurium (Te) with open-boundary conditions along [010] direction. The black solid lines $\sim$100 cm-1 correspond to obstructed atomic-insulator-like phonon modes, arising from dangling-bond vibrations at the surface. (b) Experimental configuration for circularly polarized Raman spectroscopy in backscattering geometry. (c) Raman spectrum of Te showing an additional surface (S) mode. The inset displays the fitting results, confirming the S-mode’s origin from surface dangling-bond vibrations. Adapted from Ref. Zhang et al. (2023b).
@@
Figure 30 (c) presents the Raman shifts of Te under circularly polarized incident/scattered light. In addition to the three Raman-active modes ($E^{1}$, $E^{2}$, and $A_{1}$), an extra S mode appears. The frequencies of all four modes align well with first-principles calculations (inset), confirming the S mode as the obstructed surface phonon Zhang et al. (2023b).
@@ IV.4 Topological Phonon Materials Database
The method of symmetry-based indicator theory and topological quantum chemistry (TQC) Po et al. (2017); Bradlyn et al. (2017); Elcoro et al. (2021) provides a powerful framework for identifying topological materials. Its classification hinges on the elementary band representations (EBRs) Michel and Zak (2001); Cano and Bradlyn (2021), which distinguish topologically trivial and nontrivial phases: A band structure adiabatically connected to an “atomic insulator” is trivial, whereas the absence of such a connection typically signals nontrivial topology Cano and Bradlyn (2021). This classification is rigorously formalized using symmetry-based indicators Po et al. (2017); Tang et al. (2019b, d, c); Song et al. (2018); Tang and Wan (2024); Zhang et al. (2019a); Vergniory et al. (2019a).
@@
Figure 31: Construction of database for topological phononic materials. (a) High-throughput computational screening of topological phononic nodal lines/rings/points. Adapted from Ref. Li et al. (2021c). (b) Topological phonon database for obstructed atomic band representations (OABRs) and orbital-selected OABRs (OOABR). Adapted from Ref. Xu et al. (2024a).
@@
To date, extensive high-throughput searches for topological materials, encompassing both electronic and phononic systems, have been conducted using first-principles methods Zhang et al. (2019a); Tang et al. (2019b); Vergniory et al. (2019a); Xu et al. (2020); Li et al. (2021c); Chen et al. (2021a); Xu et al. (2024a). From the symmetry perspective, phonons behave analogously to spinless electrons, with their three vibrational degrees of freedom resembling $p_{x,y,z}$ orbitals Liu et al. (2018). Consequently, the complete classification of topological phononic materials can be achieved within the framework of single-valued representations of the 230 space groups. Pioneering this effort, Li et al. (2021c) performed a high-throughput screening of topological phononic materials, identifying systems hosting phononic nodal lines and Weyl points, as shown in Fig. 31 (a). By computing phononic band structures along high-symmetry paths, they systematically classified materials with these topological features. More recently, Xu et al. (2024a) expanded this approach by screening phononic obstructed atomic bands using topological quantum chemistry, paving the way for further experimental studies, as shown in Fig. 31 (b).
V From Topological Phonons to Chiral Phonons
“Chirality”  [^1] In some references, e.g. Barron (2021), the concept of chirality is classified into two catogies: the true chirality and false chirality. The true chirality refers to the existence of two enantiomer states that are interchanged only by spatial inversion operation $\mathcal{P}$ but NOT by time-reversal operation $\mathcal{T}$; however, for the false chirality, the two enantiomer states are interchanged by both $\mathcal{P}$ and $\mathcal{T}$. The distinction between true and false chirality is essential to the problem of homochirality origin of our natural world including DNA. The homochiral phenomenon must have a true chiral origin. in phonons is multifaceted and can be interpreted diversely across various disciplines. For example, it can be defined by the Chern number in the field of topological band theory, and in this case, chiral phonons are the Weyl phonons with nonzero Chern number, which have been identified through various experimental observations Zhang et al. (2018a); Miao et al. (2018); Zhang et al. (2019b, 2020b); Li et al. (2021a); Jin et al. (2022); Zhang et al. (2023b). In another example, angular momentum (AM, $l_{\nu q}$) is introduced to describe the chiral atomic motions in the real space McLellan (1988); Zhang and Niu (2014); Hamada et al. (2018); Hamada and Murakami (2020b); Ueda et al. (2023); Zhang et al. (2024b), and the ones with nonzero AM was also referred to as chiral phonons, or circularly polarized phonons. The sign of AM also reflects the chirality of phonons, but this definition is convention-dependent because AM is a pseudo-vector. To address this issue, an alternative definition of chiral phonons based on helicity has been introduced, as discussed in Zhang et al. (2025a). Phonon helicity is a pseudo-scalar, making it convention-independent under SO(3) operations. The sign of the helicity directly indicates the chirality of circularly polarized phonons, providing a more robust and unambiguous characterization of their chiral nature. The relationship between helicity and angular momentum $l_{\nu q}$ is illustrated in Fig. 32, where $\nu$ indicates the phonon mode and $q$ is the momentum. Figure 32 shows that chiral phonons defined by the helicity can be nonchiral, even in the cases where the AM is nonzero.
According to topological band theory, the Chern number is related to the eigenvalue of (screw) rotational symmetry $\mathcal{C}_{n}$, which is also known as pseudo-angular momentum (PAM) Yao et al. (2008); Fang et al. (2012); Zhang and Niu (2015); Tsirkin et al. (2017); Zhang and Murakami (2022). That is, PAM can diagnose topological phonons. Recently, PAM also has been widely used to identify chiral phonons Zhu et al. (2018); Ishito et al. (2023c); Zhang et al. (2023b). However, Zhang et al. (2025a) highlights that there is no intrinsic connection between PAM and AM unless symmetry analysis and first-principles calculations are incorporated as supplementary tools. For instance, Zhang et al. (2023b) demonstrates that Weyl phonons and chiral phonons defined by AM/helicity are connected Zhang et al. (2020b); Li et al. (2021a); Zhang et al. (2023b), but this connection relies on a combination of topological band theory, symmetry analysis, and first-principles calculations Zhang et al. (2025b).

### v2

@@ IV.1.3 Body-centered Silicon: $\mathbb{Z}_{2}$ Dirac Phonon
Figure 25: $\mathbb{Z}_{2}$ Dirac phonon in body-centered silicon with space group $Ia\bar{3}$. (a) Crystal structure of silicon. (b) Bulk Brillouin zone (BZ) and (001) surface BZ. (c) Phonon band structure, where the colored dots denote two distinct $\mathbb{Z}_{2}$ Dirac phonons, $P_{1}P_{3}$ and $P_{2}P_{2}$, characterized by different irreducible representations (irreps). (d)-(e) Surface states and Fermi arcs associated with the $\mathbb{Z}_{2}$ Dirac phonons of irreps $P_{1}P_{3}$. (f)-(g) Surface states and Fermi arcs corresponding to the $\mathbb{Z}_{2}$ Dirac phonons of irreps $P_{2}P_{2}$. Adapted from Ref. zhang2023parallel .
As discussed in Sec. II.3, conventional Dirac phonons are topologically trivial because they consist of two Weyl phonons with opposite Chern numbers, which do not guarantee gapless helicoid surface states. To provide an intuitive understanding of $\mathbb{Z}_{2}$ Dirac phonons and their characteristic anti-parallel helical surface states, we will take an experimentally synthesized material candidate as an example in this subsection, showing their topological properties zhang2023parallel .
@@
Figure 25 (d) displays the (001) surface states for the $P_{1}P_{3}$ Dirac points identified in Fig. 25 (c), with the corresponding $k$-path shown in the surface arc calculation. The (001) surface Brillouin zone possesses two $\mathcal{TG}$ symmetries, where $(\mathcal{TG}_{x})^{2}=-1$ and $(\mathcal{TG}_{y})^{2}=-1$ along the BZ boundaries $\bar{M}$$\bar{X}$ and $\bar{M}$$\bar{Y}$. These symmetries enforce Kramers-like degeneracy, resulting in doubly degenerate surface states along these boundaries. The $\mathbb{Z}_{2}$ Dirac phonons at $P$ and $P^{\prime}$ project onto the BZ corner ($\bar{M}$), generating two pairs of anti-parallel surface states along the path $\bar{Y}$$\bar{X}$$\bar{Y}^{\prime}$$\bar{X}^{\prime}$$\bar{Y}$ in Fig. 25 (d). Four anti-parallel helical surface states cross along $\bar{X}$$\bar{M}$$\bar{X}^{\prime}$, with their gapless intersection protected by $\mathcal{TG}_{x}$. All topological surface states ultimately merge into the $\mathbb{Z}_{2}$ Dirac cone at $\bar{M}$.
Similarly, Figs. 25 (f) and (g) show the surface states and arcs for the $P_{2}P_{2}$ Dirac phonon, which also exhibit two pairs of anti-parallel helical surface states around $\bar{M}$. Thus, both the $P_{2}P_{2}$ and $P_{1}P_{3}$ Dirac phonons in silicon’s phonon spectrum are $\mathbb{Z}_{2}$ Dirac phonons, each hosting two pairs of anti-parallel helical surface states (termed anti-parallel quad-helical surface states in zhang2023parallel ). The (anti-)parallel quad-helical surface states represent the maximal number of surface states achievable for (Dirac) Weyl phonons in solids, as constrained by onsite and crystalline symmetries.
Figure 26: (a) Conventional and primitive unit cells of MoB2. (b) Phonon spectra of MoB2. (c) Topological nodal line identification in MoB2, determined through compatibility relations and symmetry-based indicator analysis. (d)-(e) Side view and [111]-oriented top view of the nodal lines near 30 meV in MoB2, exhibiting a quantized $\pi$ Berry phase. (f) Bulk Brillouin zone and [010]-projected surface BZ of MoB2. (010) surface local density of states (LDOS), showing both bulk-derived modes and surface-unique modes (left panel: integrated intensity). Drumhead surface phonon states (within white dashed rectangles) connecting nodal line crossings display maximal LDOS, as evidenced by sharp intensity peaks at nodal-line energies (shaded regions). Adapted from Ref. zhang2019phononic .
@@ IV.2 Topological Phonons with 1D Degeneracies
For MoB2, our symmetry analysis yields the indicator $z_{2,1}z_{2,2}z_{2,3}z_{2,4}z_{4}=(1110)$. Following the theory outlined in Sec. II.4 (Eqs. 80 and 81) song2018diagnosis ; zhang2020diagnosis , this indicator predicts:
@@
To demonstrate the topological surface states associated with these $\mathcal{PT}$-protected nodal lines, we present the local density of states (LDOS) along the (010) direction in Fig. 26 (f), corresponding to the surface Brillouin zone shown in Fig. 26 (e). Unlike Weyl points that exhibit helicoid surface modes fang2016topological ; zhang2018double ; miao2018observation ; zhang2020twofold ; zhang2022z ; zhang2023parallel , the nodal lines generate characteristic drumhead-shaped surface states, as indicated by the white dashed box. Notably, while the bulk phonon density of states shows minima near $\sim$7.25 THz and $\sim$23 THz, the surface modes display a pronounced peak in LDOS at these energies (Fig. 26 b). The flat surface states could induce electronic anomalies through electron-phonon coupling, potentially leading to unique surface phenomena.
@@
In 2D systems, Eq. (83) typically defines isolated $\bm{k}$-points, corresponding to 2D Dirac semimetals Li2020Feb ; Jin2018Dec . In 3D systems, however, the same condition can be satisfied along closed $\bm{k}$-space contours, giving rise to topological phononic nodal lines. Numerous material candidates hosting such nodal lines have been theoretically predicted and experimentally observed. Based on their geometric configurations, these nodal lines can be categorized into several distinct types: helical nodal lines zhang2019phononic , straight nodal lines Li2020Jan ; Liu2021Jul , nodal rings Jin2018PRB , nodal chains Chen2021Oct , and nodal net Liu2021Jul_CuCl . Phononic topological nodal lines exist ubiquitously in centrosymmetric solids liu2022ubiquitous .
Straight nodal lines along high-symmetry lines (HSLs) with $C_{n}$ rotational symmetry can be categorized into two fundamental types based on their transverse dispersion: linear and quadratic types Liu2021Jul_CuCl ; liu2022ubiquitous ; qin2024diverse . This classification applies particularly to systems with threefold rotational symmetry ($n=3$), where the low-energy effective Hamiltonian takes the general form:
@@
where $k_{\pm}=k_{1}\pm ik_{2}$ with $k_{1}$ and $k_{2}$ being the wave vector perpendicular to the rotational axis. $A$ and $B$ are complex constants that do not depend on $k_{1}$ and $k_{2}$. The location of nodal line is determined by $Ak_{-}+Bk^{2}_{+}=0$ which give rise to 4 nodal lines: One located at the rotation axis along $k_{3}$ with Berry phase $\pi$, while the other three located at generic positions which are related by $C_{3}$ rotation and have Berry phase $\pi$, i.e.,
$\begin{split}(0,0,k_{3}):&\textrm{Berry phase}\penalty 10000\ +\pi\\ \left(k_{0}\cos\frac{\theta_{0}}{3},k_{0}\sin\frac{\theta_{0}}{3},k_{3}\right):&\textrm{Berry phase}\penalty 10000\ -\pi\\ \left(k_{0}\cos\frac{\theta_{0}+2\pi}{3},k_{0}\sin\frac{\theta_{0}+2\pi}{3},k_{3}\right):&\textrm{Berry phase}\penalty 10000\ -\pi\\ \left(k_{0}\cos\frac{\theta_{0}-2\pi}{3},k_{0}\sin\frac{\theta_{0}-2\pi}{3},k_{3}\right):&\textrm{Berry phase}\penalty 10000\ -\pi\end{split}$ (85)
where $k_{0}=|-\frac{A}{B}|$ and $\theta_{0}=\mathrm{arg}\left(-\frac{A}{B}\right)$. Nodal lines can be classified into two fundamental types based on their transverse dispersion: (i) linear nodal lines with $\pi$ Berry phase (topologically nontrivial) and (ii) quadratic nodal lines with $2\pi$ Berry phase liu2022ubiquitous ; qin2024diverse . The topological character is determined by the dominant term in the effective Hamiltonian: when the linear term dominates ($|A/B|\gg 1$), the system exhibits a $\pi$ Berry phase, whereas dominance of the quadratic term ($|A/B|\ll 1$) yields a $2\pi$ Berry phase. For systems with $C_{4}$ or $C_{6}$ symmetry, the linear term in Eq. (84) vanishes, allowing only a single straight nodal line along the rotational axis. In these cases, quadratic nodal lines emerge with Berry phase $\pm 2\pi$, corresponding to a winding number $n_{W}=\pm 2$ liu2022ubiquitous . This distinct topological signature arises from the doubled winding of the pseudospin texture around the nodal line.
Figure 27: Two types of unconventionality. Atomic vibrations act as $p$ orbitals. (a) Phonon modes contributed by the vibration center at an empty site. (b) Phonon modes contributed by the vibration center at an atom site, yet show a non-atomic vibration orbitals (non-$p$ orbital symmetry). (c) Phonon spectra of the monolayer black phosphorous. (d) Phonon spectra for the monolayer black phosphorous with an open boundary condition, where there are floating obstructed surface phonon modes $\sim$ 9 THz. The real space vibration for the surface phonon mode marked by the green hexagon is shown in (e). Adapted from Ref. zhang2023unconventional .
@@ IV.3 Obstructed Atomic Insulator-like Topological Phonons
Phonons represent quantized collective vibrations of atomic lattices, where each mode corresponds to a distinct pattern of atomic displacements. However, certain phonon modes exhibit unconventional origins: they may arise from vibrations localized at empty atomic sites or involve non-atomic orbital symmetries (e.g., non-$p$-orbital-like vibrations), as illustrated in Figs. 27 (a)–(b). These anomalous modes, termed obstructed phonon modes song2020twisted ; peri2020experimental ; xu2021three ; gao2022unconventional ; xu2024filling ; xu2024catalog zhang2023unconventional ; zhang2023weyl ; ma2023obstructed , defy conventional descriptions of lattice dynamics. The unconventional nature of these modes can be identified through either real-space invariants or open-boundary calculations zhang2023unconventional ; zhang2023weyl . The former one quantifies the mismatch between atomic positions and vibrational centers, while the latter one reveals localized surface modes (“floating states”) detached from bulk phonon spectra. Such modes typically emerge from dangling-bond vibrations on crystal surfaces and are intimately linked to the material’s symmetry and electronic structure. Their spectral isolation and spatial localization make them promising for controlling phonon-mediated phenomena at interfaces or defects.
As a concrete example, Figs. 27 (c) and (d) display the phonon spectra of monolayer black phosphorus under periodic and open boundary conditions, respectively zhang2023unconventional . Notably, the open-boundary spectrum reveals floating obstructed surface phonon modes near 9 THz, which are absent in the periodic calculation. The real-space vibrational pattern of one such surface mode (indicated by the green hexagon in Fig. 27 (d) is shown in Fig. 27 (e), clearly demonstrating its localized character at the edge.
Figure 28: (a) First-principles calculations of right-handed tellurium (Te) with open-boundary conditions along [010] direction. The black solid lines $\sim$100 cm-1 correspond to obstructed atomic-insulator-like phonon modes, arising from dangling-bond vibrations at the surface. (b) Experimental configuration for circularly polarized Raman spectroscopy in backscattering geometry. (c) Raman spectrum of Te showing an additional surface (S) mode. The inset displays the fitting results, confirming the S-mode’s origin from surface dangling-bond vibrations. Adapted from Ref. zhang2023weyl .
@@
Figure 28 (c) presents the Raman shifts of Te under circularly polarized incident/scattered light. In addition to the three Raman-active modes ($E^{1}$, $E^{2}$, and $A_{1}$), an extra S mode appears. The frequencies of all four modes align well with first-principles calculations (inset), confirming the S mode as the obstructed surface phonon zhang2023weyl .
@@ IV.4 Topological Phonon Materials Database
The method of symmetry-based indicator theory and topological quantum chemistry (TQC) po2017symmetry ; bradlyn2017topological ; Elcoro2021Oct provides a powerful framework for identifying topological materials. Its classification hinges on the elementary band representations (EBRs) Michel2001Feb ; Cano2021Mar , which distinguish topologically trivial and nontrivial phases: A band structure adiabatically connected to an “atomic insulator” is trivial, whereas the absence of such a connection typically signals nontrivial topology Cano2021Mar . This classification is rigorously formalized using symmetry-based indicators po2017symmetry ; Tang2019Feb ; Tang2019Mar ; Tang2019May ; song2018diagnosis ; Tang2024Dec ; zhang2019catalogue ; Vergniory2019Feb .
@@
Figure 29: Construction of database for topological phononic materials. (a) High-throughput computational screening of topological phononic nodal lines/rings/points. Adapted from Ref. Li2021Feb . (b) Topological phonon database for obstructed atomic band representations (OABRs) and orbital-selected OABRs (OOABR). Adapted from Ref. Xu2024May .
@@
To date, extensive high-throughput searches for topological materials, encompassing both electronic and phononic systems, have been conducted using first-principles methods zhang2019catalogue ; Tang2019Feb ; Vergniory2019Feb ; Xu2020Oct ; Li2021Feb ; Chen2021Aug ; Xu2024May . From the symmetry perspective, phonons behave analogously to spinless electrons, with their three vibrational degrees of freedom resembling $p_{x,y,z}$ orbitals liu2018berry . Consequently, the complete classification of topological phononic materials can be achieved within the framework of single-valued representations of the 230 space groups. Pioneering this effort, Li2021Feb performed a high-throughput screening of topological phononic materials, identifying systems hosting phononic nodal lines and Weyl points, as shown in Fig. 29 (a). By computing phononic band structures along high-symmetry paths, they systematically classified materials with these topological features. More recently, Xu2024May expanded this approach by screening phononic obstructed atomic bands using topological quantum chemistry, paving the way for further experimental studies, as shown in Fig. 29 (b).
V From Topological Chirality to Rotational Chirality
In addition to topological phonons, chiral phonons have also attracted significant interest. Unlike topological phonons, which are rigorously defined by topological invariants, the notion of chirality  [^2] In some references, e.g. barron2021symmetry , the concept of chirality is classified into two catogies: the true chirality and false chirality. The true chirality refers to the existence of two enantiomer states that are interchanged only by spatial inversion operation $\mathcal{P}$ but NOT by time-reversal operation $\mathcal{T}$; however, for the false chirality, the two enantiomer states are interchanged by both $\mathcal{P}$ and $\mathcal{T}$. The distinction between true and false chirality is essential to the problem of homochirality origin of our natural world including DNA. The homochiral phenomenon must have a true chiral origin. and chiral phonons can vary across different research contexts. For example, in topolgical materials, the chirality of Weyl phonons is defined by the sign of the Chern number, corresponding to topological chirality that has been experimentally confirmed zhang2018double ; miao2018observation ; zhang2019phononic ; zhang2020twofold ; li2021observation ; jin2022chern ; zhang2023weyl . Another broadly used quantity is the phonon angular momentum (AM, $l_{\nu q}$). It characterizes the circular motion of atoms in real space and corresponds to the rotational chirality of phonons mclellan1988angular ; zhang2014angular ; hamada2018phonon ; hamada2020phonon ; ueda2023nature ; zhang2024understanding . In this respect, chiral phonons are defined as phonon modes with nonzero AM bermudez_chirality_2008 ; bermudez2008chirality ; bermudez2008hyper ; zhang2015chiral . In earlier literatures, these phonons are also called circularly polarized phonons pine1969linear ; johnson1982angular ; rebane1983faraday or axial phonons juraschek2025chiral . More recently, another new concept, the pseudo-angular momentum (PAM) that is defined on (screw) rotational symmetry $\mathcal{C}_{n}$, is introduced to the phonon community yao2008valley ; zhang2015chiral ; zhu2018observation ; HgS_chiral_2023 ; zhang2023weyl . However, as emphasized in zhang2025thechirality , there is no intrinsic one-to-one correspondence between PAM and AM unless symmetry analysis and first-principles calculations are employed as supplementary tools.